# Lazy-DaSH: Lazy Approach for Hypergraph-based Multi-robot Task and Motion Planning

### v1

@@ Abstract
We introduce Lazy-DaSH, an improvement over the recent state of the art multi-robot task and motion planning method DaSH, which scales to more than double the number of robots and objects compared to the original method and achieves an order of magnitude faster planning time when applied to a multi-manipulator object rearrangement problem. We achieve this improvement through a hierarchical approach, where a high-level task planning layer identifies planning spaces required for task completion, and motion feasibility is validated lazily only within these spaces. In contrast, DaSH precomputes the motion feasibility of all possible actions, resulting in higher costs for constructing state space representations. Lazy-DaSH maintains efficient query performance by utilizing a constraint feedback mechanism within its hierarchical structure, ensuring that motion feasibility is effectively conveyed to the query process. By maintaining smaller state space representations, our method significantly reduces both representation construction time and query time. We evaluate Lazy-DaSH in four distinct scenarios, demonstrating its scalability to increasing numbers of robots and objects, as well as its adaptability in resolving conflicts through the constraint feedback mechanism.
@@ I Introduction
Figure 1: A comparison of the search space scope during the search processes of DaSH and Lazy-DaSH. The introduction of the task query phase and lazy motion validation distinguishes our approach from DaSH, as highlighted in red. The task query narrows the search space, while lazy motion validation considers only motions in the candidate plan, reducing the computational cost of motion space construction. While both DaSH and Lazy-DaSH iteratively update representations upon plan failure, Lazy-DaSH employs a constraint feedback mechanism within a hierarchical framework to effectively manage both task-level and motion-level constraints, as illustrated in Fig. 2 and Algorithm 1.
In our prior work, we presented the Decomposable State Space Hypergraph (DaSH) framework [2], a hybrid approach for MR-TMP which seeks to focus computation effort only where coordination is needed through a more efficient hypergraph-based representation of the task space than traditional graph-based composite methods and produces smaller search spaces when considering the motion feasibility for groups of coupled robots and tasks. The original DaSH framework computes the motion feasibility within each of these smaller search spaces and then uses this information within a joint task and motion planning query (Fig. 1). This leads to very fast query times, but the construction time required to compute representations can be excessive, as high levels of task and motion coordination are required with the growing number of robots and objects.
This paper presents Lazy-DaSH, which reduces overall computational effort by lazily computing motion feasibility only within the set of decoupled planning spaces required by the current task plan. This requires the computation of a task plan before considering motion feasibility, earlier in the process than the original DaSH framework. This can be understood as selectively scoping the potentially relevant search space during the high-level planning process to reduce the computational effort required for the subsequent representation construction phase (Fig. 1). To further minimize computational overhead, motion planning is performed lazily, considering only the current task plan. If a required motion is found to be infeasible, the task planning phase incorporates this information to generate a new task plan, leading to an alternative motion plan. This allows the planner to reduce the cost of expensive motion validation (e.g., collision checking) and focus on sequencing and identifying critical tasks and motions, calling motion validation only when necessary.
@@
  • A task planning layer into the DaSH hierarchy, featuring queryable task planning representation and task query strategy.
  • Lazy motion validation of robot state spaces included in the task plan, with a task and motion constraint feedback mechanism for infeasible motions.
  • Scalable and efficient replanning, scaling to twice the number of both robots and objects and achieving an order of magnitude faster planning times than DaSH in diverse multi-manipulator object rearrangement scenarios.
@@ II-B Task and Motion Planning
Satisfaction-first approaches focus on satisfying constraints related to continuous parameters (e.g., object positions and robot configurations) before creating an action sequence [13, 14, 15]. They are efficient when it is easier to sample and test these parameters upfront rather than repeatedly attempting to fit them into an action sequence that might not work. Thus, they often involve a cycle of sampling and testing for feasibility.
@@ II-C Multi-robot Task and Motion Planning
Multi-robot task and motion planning extends the fundamental concepts of task and motion planning to complex tasks that require collaboration between multiple robots. This problem is significantly more complex when factoring in robot interactions in coordination, collision avoidance, and task allocation and scheduling [20].
@@
Multi-manipulator object rearrangement is an important problem in MR-TMP requiring complex multi-robot coordination. The three classes of approaches—coupled, decoupled, and hybrid—are applicable in this context as well. Most MR-TMP problems require a high level of coordination, and the decoupled approach often fails to find a feasible solution. On the other hand, the coupled approach is capable of finding a highly coordinated solution but is not scalable due to the size of the search space.
@@
Most research on the multi-manipulator object rearrangement problem has focused on developing efficient representations for the coupled approach. For example, the authors in [21] utilize a shared manipulator workspace to create a shared space graph, enabling reasoning about multi-robot cooperation and adapting a path planning heuristic for multi-manipulator tasks. Moreover, a series of studies [22, 23, 24] has focused on developing a concise object mode graph [15]. This graph captures valid transitions for pick, place, and hand-over actions and serves as a heuristic for guiding multi-modal motion planning [15]. Building on the work in [23], [25] utilizes the object-centric mode graph for multiple objects and applies multi-agent pathfinding techniques to generate non-conflicting sequences of object modes. For more complex assembly tasks, [26] employs a Mixed-Integer Linear Programming method for task assignment but did not effectively demonstrate scalability with respect to the number of robots, primarily due to the computational demands of roadmap generation and annotation processes.
@@
Hybrid methods, such as [20], introduced a Conflict Based Search (CBS) variant to address the MR-TMP problem with multiply decomposable tasks. While this approach demonstrated the effectiveness of hybrid methods in complex multi-robot task and pathfinding problems, it has not been validated for higher-dimensional planning scenarios, such as multi-manipulator object rearrangement problems. DaSH [2], a recent hybrid method, employs a hypergraph representation to concisely model hybrid robot state spaces and has been validated in the multi-manipulator object rearrangement problem. While it shows performance improvements—three orders of magnitude faster than the benchmark presented in [25]—its scalability is limited, particularly in scenarios with complex constraints. For example, in tasks requiring geometric constraints or multiple steps for completion, DaSH is limited in leveraging task-inferred constraint information during the query process. This limitation arises because a combined task and motion query in DaSH relies only on motion-collision constraints for replanning when the query fails, rather than incorporating task-specific constraints. In Section V, we demonstrate improved performance of Lazy-DaSH in environments with geometric constraints, even with a higher number of robots and objects.
@@ III-A Task Space
In the MR-TMP framework [2], the task space $\mathcal{T}$ is defined by three key components: the movable bodies $\mathcal{B}$, the configuration spaces $\mathcal{W}$, and the constraints $\mathcal{C}$. Each task space element $T_{i}=(B_{i},W_{i},C_{i})\in\mathcal{T}$ consists of a subset of the movable bodies $B_{i}\subseteq\mathcal{B}$, its associated configuration space $W_{i}\subseteq\mathcal{W}$, and a set of constraints $C_{i}\subseteq\mathcal{C}$ that determine the validity of configurations within $W_{i}$. An MR-TMP problem has a set of admissible task space elements $\mathcal{T}^{*}\subseteq\mathcal{T}$, where the movable bodies and constraints of $T_{i}\in\mathcal{T}^{*}$ define a valid set of configurations for the entire system. The planner can consider transitions between sets of task space elements. These may reflect changes in the compositions of the subsets of moveable bodies and/or in the constraint sets included in elements.
@@ III-B Multi-manipulator Object Rearrangement Problem
In scenarios where manipulators are used to rearrange objects, the admissible task space elements $\mathcal{T}^{*}\subseteq\mathcal{T}_{\texttt{MANIP}}$ include robots, objects, and every possible combination of manipulator-object grasp pairs. These task space elements encompass not only the manipulators and objects themselves but also the manipulator-object pairs, along with their associated configuration spaces and constraints. For instance, task space elements for objects are characterized by the workspace’s configuration space, constrained to ensure the object is on a stable surface. For manipulators, the elements are described by the robot’s joint configuration space, with constraints to prevent self-collision. The elements for manipulator-object pairs are defined by the composite configuration space of both the robot and the object, constrained by the condition that the end-effector securely grasps the object.
Transitions are performed by pick, place, or hand-over actions. A pick action on an object, for example, removes the constraint that the object is on a stable surface and adds a constraint that the object is in a stable grasp by the robot while changing the composition from independent elements, one with a robot and one with an object, to a coupled robot-object composition in the post-transition element. Transitions between task space elements are only possible at valid configurations between objects and robots for grasping. Consequently, a solution to this problem involves a discrete sequence of task space elements that specify which robot grasps which object, along with continuous motion paths within the configuration space of each task space element.
Figure 2: Illustration of the hierarchical structure of the proposed Lazy-DaSH, showing two manipulators ($R_{1}$ and $R_{2}$) rearranging an object ($O_{1}$). The task query phase and the task constraint feedback scheme, which distinguish it from DaSH, are highlighted with red lines. This feature is also emphasized in Algorithm 1. Each layer of the hierarchy is detailed in corresponding sections.
@@ IV The Lazy-DaSH Method
The overall structure of Lazy-DaSH is illustrated in Fig. 2. As in DaSH [2], we utilize a hierarchical representation structure, using hypergraph to represent the task spaces and motion spaces. As oppose to a standard graph, a (directed) hypergraph utilizes hyperedges (or hyperarcs) that connect multiple vertices simultaneously [27]. This is particularly useful for compactly encoding inter-robot interactions and constraints within a relatively small size compared to a standard graph [2].
In contrast to DaSH, Lazy-DaSH employs a two-stage hierarchical query approach: a task query within the task space hypergraph followed by a motion query within the motion hypergraph. This creates a strict hierarchical structure with a task planning layer and motion planning layer, where each layer has its corresponding representation construction phase and its query phase. The representation construction phase encodes the satisfaction of task and motion constraints between robots and objects, while the query phase sequences the constraint-satisfied task and motion.

### v2

@@ Abstract
We introduce Lazy-DaSH, an improvement over the recent state of the art multi-robot task and motion planning method DaSH, which scales to more than twice the number of robots and objects while achieving an order of magnitude faster planning when applied to a multi-manipulator object rearrangement problem. We achieve this improvement through a hierarchical approach, where a high-level task planning layer identifies planning spaces required for task completion, and motion feasibility is validated lazily only within these spaces. In contrast, DaSH precomputes the motion feasibility of all possible actions, resulting in higher costs for constructing state space representations. Lazy-DaSH ensures efficient query performance by utilizing a hierarchical constraint feedback mechanism, effectively conveying motion feasibility to the query process while incrementally expanding the task and motion space representations when failures are detected, so the search space grows only as needed. By maintaining smaller state space representations, our method significantly reduces both representation construction time and query time. We evaluate Lazy-DaSH in four scenarios, demonstrating its scalability with increasing numbers of robots and objects, as well as its adaptability in resolving conflicts through the constraint feedback.
@@ I Introduction
Figure 1: A comparison of the search space scope during the search processes of DaSH and Lazy-DaSH. The introduction of the task space expansion, task query, task constraint detection, task and motion constraint feedback, and lazy motion validation phases distinguishes our approach from DaSH, as highlighted in red. The task query narrows the search space, while lazy motion validation considers only motions in the candidate plan, reducing the computational cost of motion space construction. The task and motion constraint feedback initiates the expansion of the task space and motion space, thereby broadening the search space in the respective planning representations. While both DaSH and Lazy-DaSH iteratively update representations upon plan failure, Lazy-DaSH employs a constraint feedback mechanism within a hierarchical framework to effectively manage both task-level and motion-level constraints, as illustrated in Fig. 2 and Algorithm 1.
The Decomposable State Space Hypergraph (DaSH) framework [2] is a hybrid approach for MR-TMP which seeks to focus computation effort only where coordination is needed through a more efficient hypergraph-based representation of the task space than traditional graph-based composite methods and produces smaller search spaces when considering the motion feasibility for groups of coupled robots and tasks. The original DaSH framework computes the motion feasibility within each of these smaller search spaces and then uses this information within a combined task and motion planning query (Fig. 1). This leads to fast query times, but the construction time required to compute representations can be excessive, as high levels of task and motion coordination are required with the growing number of robots and objects. In addition, queries that rely solely on motion-level representations often struggle in environments with task-oriented constraints.
This paper presents Lazy-DaSH, which extends DaSH by managing representation size within a hierarchical structure. Instead of exhaustively expanding the motion-level representation, Lazy-DaSH selectively expands both task- and motion-level representations under the guidance of task and motion constraint management. This selective expansion lowers the cost of representation construction, while motion feasibility is evaluated lazily, only within the decoupled planning spaces required by the current task plan. By deferring motion validation and using constraint management to refine the search space, Lazy-DaSH reduces the overhead of operations such as collision checking and focuses computational effort on sequencing and identifying critical tasks and motions. Through the integration of efficient representation control, hierarchical constraint management, and lazy motion validation, Lazy-DaSH achieves a balance between computational efficiency and effective reasoning over task-oriented constraints during the query process.
@@
  • A new task planning layer within the DaSH hierarchy, equipped with a queryable task representation and task query strategy.
  • Lazy motion validation of robot state spaces included in the task plan, improving computational efficiency.
  • Hierarchical constraint management for efficient control of representation size, expanding task and motion space representations only when necessary.
  • Scalable and efficient replanning, handling more than twice the number of robots and objects compared to DaSH and achieving up to two orders of magnitude faster planning times in multi-manipulator object rearrangement scenarios.
@@ II-A Motion Planning
The main computational cost in this process comes from validating sampled configurations and edges, which requires repeated collision checking against the environment [9, 10]. To address this, methods such as lazy evaluation of edges [11, 5], reuse of previous collision-check results [12], adaptive or biased sampling strategies [13], and hybrid search-plus-optimization planners  [14, 15] have been proposed to reduce unnecessary checks while still ensuring correctness.
@@ II-B Task and Motion Planning
Satisfaction-first approaches focus on satisfying constraints related to continuous parameters (e.g., object positions and robot configurations) before creating an action sequence [20, 21, 22]. They are efficient when it is easier to sample and test these parameters upfront rather than repeatedly attempting to fit them into an action sequence that might not work. Thus, they often involve a cycle of sampling and testing for feasibility check.
@@
In all of these methods, the key consideration is identifying which constraints must be captured and how they should be managed to establish valid plans. The constraints that determine task or motion feasibility are often geometric or kinematic in nature and can be represented through task-space abstractions such as scene graphs [27], symbolic-geometric models [28], or constraint-based formulations [29]. By efficiently encoding these constraints across the hierarchical structure, planners can integrate motion-level feedback (e.g., collision checks or parameter failures) into task-level reasoning, thereby refining plans and guiding queries more effectively [16, 28].
@@ II-C Multi-robot Task and Motion Planning
Multi-robot task and motion planning extends the fundamental concepts of task and motion planning to complex tasks that require collaboration between multiple robots. This problem is significantly more complex when factoring in robot interactions in coordination, collision avoidance, and task allocation and scheduling [30]. The primary challenge for multi-robot planning is maintaining the necessary coordination while accounting for the combinatorial growth of the search space. There are three standard approaches methods take to address this challenge: decoupled, which sacrifices coordination for smaller search spaces by planning for robots individually; coupled, which accepts the size of the search space in order to maintain coordination; and hybrid, which attempts to leverage the strengths of both coupled and decoupled while minimizing their weaknesses.
@@
Most MR-TMP problems require a high level of coordination, and the decoupled approach often fails to find a feasible solution. On the other hand, the coupled approach is capable of finding a highly coordinated solution but is not scalable due to the size of the search space. This creates a fundamental trade-off between solution optimality and computational scalability, which has driven the development of various planning strategies.
At one end of the MR-TMP spectrum, coupled approaches establish a benchmark for solution quality by aiming for theoretical guarantees like optimality. For example, the method proposed in [31] formalizes the multi-robot planning problem and provide asymptotically optimal baseline planners by operating in the full composite configuration space of all robots with multiple target goals. Such methods provide a foundational standard for asynchronous path quality but assumes the given task assignment.
To make planning tractable for large-scale, long-horizon applications, other approaches prioritize computational scalability [32]. Decoupled methods achieve this by planning for single or subgroup of robots individually and then coordinating their actions in a post-processing step to resolve conflicts. For example, [33] presents breaking down a large problem into a series of smaller, more manageable subproblems to solve the long-horizon planning problem. The authors present a scalable planner for complex construction tasks by strategically decomposing the global problem into a sequence of smaller optimization problems. This approach enables the coordination of large, heterogeneous teams for tasks that are beyond the scope of fully coupled methods, trading theoretical optimality for the ability to solve highly complex, real-world scenarios.
Another key consideration for optimality and computational efficiency in MR-TMP is the execution model, which can be either synchronous or asynchronous. Synchronous approaches, where robots operate in lockstep, are often inefficient as faster robots are forced to wait for slower ones [34]. Composite planning often enables asynchronousity by exposing partial orders over actions and scheduling them to compact the overall makespan, but computing globally optimal asynchronous schedules is typically intractable at scale [35]. To deal with the computational intractability of asynchronous planning, [35] first computes a Temporal Plan Graph (TPG), and post-process a completed plan into a more flexible partial-order representation, to make the system robust to real-world delays. In addition, asynchronous planning can be done by factoring time into the planner itself to create asynchronous motion among the robots involved in the tasks [33, 36].
Hybrid methods aim to balance the trade-offs between coupled and decoupled approaches. Many of these methods adapt the two-level structure of the grid world multi-agent pathfinding method, Conflict-Based Search (CBS), for the more complex TAMP domain. For example, TMP-CBS [30] maps the CBS methodology to task planning to handle subtask dependencies. This approach decouples the problem at its low-level search, where it plans optimal paths for individual robots. The high-level search, however, operates in a coupled manner by identifying conflicts between these individual plans and imposing constraints to resolve them. While this approach demonstrated the effectiveness of hybrid methods in complex multi-robot task and pathfinding problems, it has not been validated for higher-dimensional planning scenarios, such as multi-manipulator task and motion planning problem. DaSH [2] employs a hypergraph representation to concisely model state spaces and queries plans in a decoupled manner and coupling tasks only when constraints require synchronization, while enabling asynchrounous planning. While powerful, its scalability is limited by its representation construction, which pre-computes motion feasibility for all possible actions. This high upfront cost becomes a bottleneck in scenarios with many robots, objects, and complex geometric constraints, as the number of potential actions grows exponentially.
II-D Multi-manipulator Object Rearrangement Problem
Multi-manipulator object rearrangement is an important problem in MR-TMP requiring complex multi-robot coordination. Rather than focusing on symbolic Planning Domain Definition Language search [37, 16], the multi-manipulator rearrangement problem emphasizes multi-modal reasoning over discrete modes (pick, place, handover, support) interleaved with continuous, collision-free motions [2, 22, 38, 39].
@@ II-D Multi-manipulator Object Rearrangement Problem
Most research on the multi-manipulator object rearrangement problem has focused on developing efficient representations to encode the multi-modality using the coupled approach. For example, the authors in [40] utilize a shared manipulator workspace to create a shared space graph, enabling reasoning about multi-robot cooperation and adapting a path planning heuristic for multi-manipulator tasks. Moreover, a series of studies [41, 42, 43] has focused on developing a concise object mode graph [22]. This graph captures valid transitions for pick, place, and hand-over actions and serves as a heuristic for guiding multi-modal motion planning [22]. Building on the work in [42], [34] utilizes the object-centric mode graph for multiple objects and applies multi-agent pathfinding techniques to generate non-conflicting sequences of object modes. For more complex assembly tasks, [44] employs a Mixed-Integer Linear Programming method for task assignment but did not effectively demonstrate scalability with respect to the number of robots, primarily due to the computational demands of roadmap generation and annotation processes.
Although representations that encode object–robot relations at both the task and motion levels provide a reasonable foundation for efficient querying, the composite state space often limits scalability when handling multiple robots and objects. It also tends to enforce synchronous planning, requiring robots to start and finish at the same global time. These restrictions reduce efficiency and frequently necessitate additional post-processing to yield a more compact task plan.
@@
DaSH [2], a recent hybrid and asynchronous MR-TMP method, employs a hypergraph representation to concisely model hybrid robot state spaces and has been validated in the multi-manipulator object rearrangement problem. It queries plans in a decoupled and parallel manner, while coupling decomposed tasks when constraints require synchronization to resolve conflicts among the decoupled tasks. While it shows performance improvements–three orders of magnitude faster than the benchmark presented in [34]–its scalability is limited, particularly in scenarios with complex constraints. For example, in tasks requiring geometric constraints or multiple steps for completion, DaSH cannot fully leverage task-inferred constraint information during the query process. This limitation arises because a combined task and motion query in DaSH relies only on motion-collision constraints for replanning when the query fails, rather than incorporating task-specific constraints. In Section V, we demonstrate improved performance of Lazy-DaSH in environments with geometric constraints, even with a higher number of robots and objects.
@@ III-A Task Space
In the MR-TMP framework [2], the task space $\mathcal{T}$ is defined by three key components: the movable bodies $\mathcal{B}$, the configuration space $\mathcal{W}$, and the constraints $\mathcal{C}$. Each task space element $T_{i}=(B_{i},W_{i},C_{i})\in\mathcal{T}$ consists of a subset of the movable bodies $B_{i}\subseteq\mathcal{B}$, its associated configuration space $W_{i}\subseteq\mathcal{W}$, and a set of task space constraints $C_{i}\subseteq\mathcal{C}$ that define the valid of configurations space, $W_{i}$, for the bodies $B_{i}$. Constraints encode the system’s physical limitations (e.g., kinematics, collisions) and task requirements (e.g., stable grasps).
An MR-TMP problem has a set of admissible task space elements $\mathcal{T}^{*}\subseteq\mathcal{T}$, where the movable bodies and constraints of $T_{i}\in\mathcal{T}^{*}$ define a valid set of configurations for the entire system. The planner can consider transitions between sets of task space elements. These may reflect changes in the compositions of the subsets of moveable bodies and/or in the constraint sets included in elements.
@@ III-B Multi-manipulator Object Rearrangement Problem
In scenarios where manipulators are used to rearrange objects, the admissible task space elements $\mathcal{T}^{*}\subseteq\mathcal{T}_{\texttt{MANIP}}$ include robots, objects, and every possible combination of manipulator-object grasp pairs. These task space elements encompass not only the manipulators and objects themselves but also the manipulator-object pairs, along with their associated configuration spaces and constraints. We define three primary task space constraints. The stability constraint $s_{c}(o_{j},q)$ for an object $o_{j}$ is satisfied when the object is placed in a stable pose at $q$. The hand-free constraint $h_{c}(r_{i})$ for a robot $r_{i}$ is satisfied when the robot is empty-handed. The formation constraint $f_{c}(r_{i},o_{j},t)$ for a robot–object pair $(r_{i},o_{j})$ is satisfied when the pair maintains a predefined stable grasp formation, where $t\in SE(3)$. Using these definitions, a task space element corresponding to a free object $o_{j}$ must satisfy $s_{c}(o_{j},q)$, while an element representing a grasped pair $(r_{i},o_{j})$ must satisfy $f_{c}(r_{i},o_{j},t)$ for both the robot and object, with $h_{c}(r_{i})$ no longer valid.
Transitions between task space elements are realized through pick, place, or hand-over actions that explicitly modify the active constraint sets, thereby moving the system to new task space elements. Each action is characterized by preconditions and postconditions defining the resulting element(s). For example, a pick action on object $o_{j}$ by robot $r_{i}$ transitions the system from a state represented by two independent task space elements–one for the free robot and one for the stable object–into a composite element representing the pair $(r_{i},o_{j})$. The preconditions for this transition require that the object is in a stable placement at configuration $q$ (i.e., $s_{c}(o_{j},q)$ holds) and that the robot can achieve a valid grasp configuration. In the resulting composite element (with both the robot and the object), the original $s_{c}(o_{j},q)$ constraint is replaced by the $f_{c}(r_{i},o_{j},t)$ constraint, capturing the stable grasp. Consequently, a solution to this problem involves a discrete sequence of task space elements, along with continuous motion paths within the configuration space of each task space element.
Figure 2: Illustration of the hierarchical structure of the proposed Lazy-DaSH, showing two manipulators ($R_{1}$ and $R_{2}$) rearranging an object ($O_{1}$). The task query phase and the task constraint feedback scheme, which distinguish it from DaSH, are highlighted with red lines. This feature is also emphasized in Algorithm 1. Each layer of the hierarchy is detailed in corresponding sections. Note that the different types of grasp modes are omitted from these figures to improve clarity of visualization and conceptual explanation.
@@ IV The Lazy-DaSH Method
In this section, we first provide an overview of the Lazy-DaSH method contrasting it with the original framework (Fig. 2). This is followed by detailed explanations of each layer of the hierarchical approach and a discussion on the properties of the method.
IV-A Overview
The overall structure of Lazy-DaSH is illustrated in Fig. 2. As in DaSH [2], we adopt a hierarchical representation, using a hypergraph to model both the task space and the motion space. In contrast to a standard graph, a (directed) hypergraph employs hyperedges (or hyperarcs) that connect multiple vertices simultaneously [45]. Whereas a graph-based representation captures the composite state space, a hypergraph captures a hybrid state space, in which task space elements couple only when enforced by transition constraints. This property is particularly useful for compactly encoding inter-robot interactions and constraints on task space elements within a relatively small representation compared to a standard graph [2].
Building on this hypergraph representation, both DaSH and Lazy-DaSH perform a representation-based query process. DaSH integrates task and motion queries into a single process by performing the query on the motion-level representation, whereas Lazy-DaSH differentiates itself by adopting a two-stage hierarchy: a task query in the task space hypergraph followed by a motion query in the motion hypergraph. This creates a strict hierarchical structure with a task planning layer and motion planning layer, where each layer has its corresponding representation construction phase and its query phase. The representation construction phase encodes the satisfaction of task and motion constraints between robots and objects, while the query phase sequences the constraint-satisfied task and motion.